# Memory in neural activity: long-range order without criticality

### v4

@@ B.2 Phase structure with varying noise levels
To further investigate the role of noise, we remove the noise term, $\sigma\eta_{i}$, from the resource equation, Eq. 2, and perform the same analysis as in the main text. The results, shown in Fig. 4, indicate that the outcomes remain qualitatively unchanged with or without the noise term in Eq. 2. This confirms that noise in the resource dynamics is not the primary driver of the observed LRO phase, thereby ruling out the possibility of a temporal Griffiths phase [33, 34].
@@ B.6 Synchronization transition
$K=\frac{1}{N}\Big{\langle}\Big{|}\sum_{i=1}^{N}e^{i\phi_{i}(t)}\Big{|}\Big{\rangle}$ (21)
where $N$ is the number of neurons and $\langle\cdot\rangle$ denotes a combined time and ensemble average. Here, $\phi_{i}(t)$ is the instantaneous phase of neuron $i$. A value of $K=1$ indicates full synchronization, while $K=0$ corresponds to completely desynchronized dynamics.
In our system, the phase $\phi_{i}(t)$ is extracted via the Hilbert transform of the neural activity $\rho_{i}(t)$:
$\tilde{\rho}_{i}^{\text{real}}(t)=$ $\rho_{i}(t)-\langle\rho_{i}\rangle,$ (22)
$\tilde{\rho}_{i}^{\text{imag}}(t)=$ $\mathcal{H}[\tilde{\rho}_{i}^{\text{real}}(t)],$
$\phi_{i}(t)=$ $\arctan\bigg{(}\frac{\tilde{\rho}_{i}^{\text{imag}}(t)}{\tilde{\rho}_{i}^{\text{real}}(t)}\bigg{)}.$
@@
Fig. 9 shows the computed Kuramoto index as a function of $\tau_{D}$. Two sharp transitions are evident at $\tau_{D}=20$ and $\tau_{D}=82$. Within the range $20\leq\tau_{D}\leq 82$, $K$ remains close to 1, yet strictly below it, indicating a moderately high degree of synchronization. This range coincides with the LRO phase identified in the main text. The elevated $K$ arises from avalanches of different sizes: long-range interactions and system-spanning events tend to promote synchronization, while smaller avalanches prevent perfect locking, keeping $K$ below 1.

### v5

@@ B.2 Phase Structure with Varying Noise Levels
To further investigate the role of noise, we remove the noise term, $\sigma\zeta_{\vec{x}}$, from the resource equation, Eq. 2, and perform the same analysis as in the main text. The results, shown in Fig. 4, indicate that the outcomes remain qualitatively unchanged with or without the noise term in Eq. 2. This confirms that noise in the resource dynamics is not the primary driver of the observed LRO phase, thereby ruling out the possibility of a temporal Griffiths phase [33, 34].
@@ B.6 Synchronization Transition
$K=\frac{1}{N}\Big\langle\Big|\sum_{\vec{x}}e^{i\phi_{\vec{x}}(t)}\Big|\Big\rangle$ (21)
where $N$ is the number of neurons and $\langle\cdot\rangle$ denotes a combined time and ensemble average. Here, $\phi_{\vec{x}}(t)$ is the instantaneous phase of the neuron at site $\vec{x}$. A value of $K=1$ indicates full synchronization, while $K=0$ corresponds to completely desynchronized dynamics.
In our system, the phase $\phi_{\vec{x}}(t)$ is extracted via the Hilbert transform of the neural activity $\rho_{\vec{x}}(t)$:
$\tilde{\rho}_{\vec{x}}^{\text{real}}(t)=$ $\rho_{\vec{x}}(t)-\langle\rho_{\vec{x}}\rangle,$ (22)
$\tilde{\rho}_{\vec{x}}^{\text{imag}}(t)=$ $\mathcal{H}[\tilde{\rho}_{\vec{x}}^{\text{real}}(t)],$
$\phi_{\vec{x}}(t)=$ $\arctan\bigg(\frac{\tilde{\rho}_{\vec{x}}^{\text{imag}}(t)}{\tilde{\rho}_{\vec{x}}^{\text{real}}(t)}\bigg).$
@@
Fig. 9 shows the computed Kuramoto index as a function of $\tau_{D}$. Two sharp transitions are evident at $\tau_{D}=20$ and $\tau_{D}=82$. Within the range $20\leq\tau_{D}\leq 82$, $K$ remains close to 1, yet strictly below it, indicating strong global phase coherence without perfect synchrony. This range coincides with the LRO phase identified in the main text.
The fact that $K$ is close to 1 reflects the fact that neural activity across the lattice is approximately periodic and broadly aligned in phase: most sites fire in a repeating, lattice-wide rhythm. However, $K$ does not reach 1 because the firing times are not exactly identical across all sites. Instead, there is a small but spatially structured timing offset from site to site within each oscillation cycle. These timing offsets are large enough to break each cycle into multiple spatiotemporal activation clusters (the avalanches we analyze), rather than a single perfectly simultaneous burst, but still small compared to the overall oscillation period. As a result, the system exhibits both (i) near-synchronous oscillations at the global level and (ii) scale-free avalanche statistics at the level of individual co-activation events.